# Multi-Agent Target Assignment and Path Finding for Intelligent Warehouse: A Cooperative Multi-Agent Deep Reinforcement Learning Perspective

### v1

@@ Abstract
Multi-agent target assignment and path planning (TAPF) are two key problems in intelligent warehouse. However, most literature only addresses one of these two problems separately. In this study, we propose a method to simultaneously solve target assignment and path planning from a perspective of cooperative multi-agent deep reinforcement learning (RL). To the best of our knowledge, this is the first work to model the TAPF problem for intelligent warehouse to cooperative multi-agent deep RL, and the first to simultaneously address TAPF based on multi-agent deep RL. Furthermore, previous literature rarely considers the physical dynamics of agents. In this study, the physical dynamics of the agents is considered. Experimental results show that our method performs well in various task settings, which means that the target assignment is solved reasonably well and the planned path is almost shortest. Moreover, our method is more time-efficient than baselines.
@@ 1 Introduction
With the rapid development of logistics related industries, new opportunities and challenges have been proposed for intelligent warehousing [1, 2]. Traditional warehouse technologies use conveyor belts and other equipment to complete material handling. They are inflexible and not easily extensible. Intelligent warehousing utilizes a multi-agent system to deliver goods to designated locations, greatly improving warehousing efficiency. Intelligent warehouse system is divided mainly into the order fulfillment system, such as the kiva system [3] and the logistics sorting center [4]. In order fulfillment systems, mobile agents move inventory pods to inventory stations. Then, workers take goods from inventory pods, and agents move the inventory pods to their original locations. In logistics sorting centers, agents take goods from loading stations and deliver them to chutes in the center of warehouse. The approach proposed in this study is based on the logistics sorting center.
Task assignment and path finding (TAPF) are two important processes in intelligent warehousing. The system first assigns a specified task to each agent according to the order requirements. Then the agent transports the goods from the origin to the destination and ensures that the path does not conflict with other agents. The task assignment and path planning problem is typically NP-hard [5], which has a large search space. Thus, directly solving this problem is difficult. In general, two steps are required to solve the TAPF problem. The first step is multi-agent task assignment (MATA), assigning tasks to agents without considering possible path conflicts between agents. The second step is path planning for all agents by using multi-agent path finding (MAPF) algorithms. Numerous studies in the literature study MATA and MAPF; we describe these methods in detail in the next section. Although solving TAPF separately reduces the difficulty of total problem, this type of method ignores the mutual influence between task assignment and path planning. Reasonable task assignment can not only effectively reduce the path length of warehouse agents and improve the operation efficiency, but also help avoid path conflicts between different agents. Therefore, it is necessary to solve the TAPF problems together. In TAPF problem, we assume that each idle agent and inventory pods are homogeneous. Thus, we can assign any task to any agent.
Recently, deep reinforcement learning (RL) has received great attention due to its remarkable successes in widely challenging domains, such as robotic control [6], multi-agent problems [7], and video games [8]. In this study, we introduce cooperative multi-agent deep RL to solve the TAPF problem to simultaneously address the TAPF problem.
@@
    To our best knowledge, this is the first work to model TAPF problem for intelligent warehouse to cooperative multi-agent deep RL, and the first to simultaneously address TAPF based on multi-agent deep RL.
@@
    In this study, the physical dynamics of the agents is considered in the path planning phase.
@@
    Experimental results show that our method performs well in various of task settings, which means the target assignment is solved reasonably and the planned path is almost shortest. Furthermore, our method is more time-efficient than the baselines.
The rest of this study is organized as follows. Section 2 discusses related work of TAPF in an intelligent warehouse. Section 3 introduces the preliminaries of the Markov Decision Process, RL, the cooperative multi-agent deep RL and the definition of MAPF and TAPF problem. Section 4 detailedly describes our method. Section 5 provides experimental results to verify the improved efficiency of our method. Section 6 presents conclusions and future work.
@@ 2.1 Multi-Agent Task Assignment
Through multi-agent task assignment algorithms, we can maximize the utilization of warehouse resources [9]. At present, multi-agent task assignment algorithms in intelligent warehouse can be classed into centralized and distributed classes according to the management mode. In the centralized task assignment class, a central control system is set up, which is responsible for task assignment, assigning tasks assigned to agents for execution [10]. Classical centralized task assignment algorithms include: Hungarian algorithm [11], tabu search algorithm [12], genetic algorithm [10], etc.. In the distributed assignment class, each agent in the warehouse plans its own task sequence according to tasks and environmental information [13]. This class method effectively reduces the load of the central control system and is more flexible and adaptable, but it may not find the global optimal solution [14]. Furthermore, the distributed assignment class mainly includes: learning-based methods [15] and market auction methods [16].
@@ 2.2 Multi-Agent Path Finding
In recent years, MAPF has become a hot research direction in the fields of computer science and robotics. Classical MAPF algorithms can be divided into optimal and sub-optimal types according to whether the solution results meet the optimality. The optimal typical MAPF algorithms include A*-based [17], conflict-based search [18], increasing cost tree search based [19], and compilation-based [20] methods. Sub-optimal classical MAPF algorithms include search-based [21], rule-based [22], compilation-based [23], bounded conflict-based search [24], and genetic algorithm-based [25] methods. However, traditional MAPF algorithms have poor real-time performance. Therefore, numerous researchers are beginning to study MAPF based on deep RL [26] to solve this problem. PRIMAL [27] is a typical MAPF algorithm based on deep RL. However, PRIMAL still solves the multi-agent problem by single-agent deep RL method [27]. In this study, we propose a method to simultaneously address target assignment and path planning from a perspective of cooperative multi-agent deep RL. Furthermore, in the simulation environment, the above-mentioned various MAPF algorithms rarely consider the physical dynamics of agents. In this study, the physical dynamics factors of the agent are considered.
@@ 2.3 Cooperative Multi-Agent Deep Reinforcement Learning
Cooperative multi-agent RL deals with systems consisting of several agents which interact in a common environment. Cooperative multi-agent RL aims to learn a policy that can make all agents work together to complete a common goal. Multi-agent deep RL has received great attention due to its capability in allowing agents to learn to make decisions in multi-agent environments through interactions. Independent Q-learning (IQL) [28] solves the multi-agent RL problem by decomposing it into a collection of simultaneous single-agent RL problems that share the same environment. However, IQL cannot solve the nonstationarity [29] problem caused by the changing policies of other agents. Value-Decomposition Networks (VDN) [30] proposes to learn to linearly decompose a team value function into a series of per-agent value functions. After that, QMIX [7] utilizes a mixing network architecture to approximate all agents’ joint state-action value. The mixing network non-linearly combines each agent’s state-action value, which is only a condition on each agent’s local observations. Furthermore, QMIX enforces the joint state action value monotonic with each agent state-action value by employing a network structural monotonicity constraint to the mixing network.
However, these above algorithms cannot handle continuous action space problems. The deep deterministic policy gradient (DDPG) [6] is a representative method in the continuous action task for a single agent. Multi-agent deep deterministic policy gradient (MADDPG) [31] extended DDPG from single agent setting to multi-agent setting by using centralised training and decentralised execution paradigm [32]. MADDPG is a representative algorithm for multi-agent continuous action problems. Thus, in this study, we model the TAPF in intelligent warehouse as a cooperative multi-agent deep RL problem and introduce MADDPG to simultaneously address this problem.
@@ 3 Preliminaries
This section provides a summary of Markov Decision Process, RL, cooperative multi-agent deep RL, and the definition of MAPF and TAPF problem.
@@ 3.3 Definition of MAPF and TAPF problem
MAPF Problem:
In the Multi-Agent Path Finding (MAPF) problem, we consider a finite-horizon Markov Decision Process defined by a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma,T)$:
  • 1.
    $\mathcal{S}$: State space.
  • 2.
    $\mathcal{A}$: Finite action space.
  • 3.
    $\mathcal{P}(s_{t+1}|s_{t},a_{t})$: State transition distribution.
  • 4.
    $r(s_{t},a_{t})$: Reward function.
  • 5.
    $\gamma$: Discount factor.
  • 6.
    $T$: Time horizon.
Each agent $n$ corresponds to an RL agent. At each time step $t$, agent $n$ selects an action $a^{n}_{t}\in\mathcal{A}$ based on its policy $\pi^{n}$. The state $s_{t}$ represents the current positions of all agents. The action $a^{n}_{t}$ can involve moving to an adjacent position or staying at the current position. The objective is to find a policy $\pi^{n}$ for each agent that maximizes the expected cumulative rewards $\mathbb{E}_{\pi,\mathcal{P}}[r_{0:T-1}]$, where $r_{0:T-1}$ is the sum of the rewards over time.
Conflicts in MAPF typically involve collisions in the selection of actions between different agents, resulting in conflicts of positions and movements.
TAPF Problem:
The Target Assignment for Multi-Agent Path Finding (TAPF) problem is defined within the framework of a finite-horizon Markov Decision Process as $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma,T)$:
  • 1.
    $\mathcal{S}$: State space.
  • 2.
    $\mathcal{A}$: Finite action space.
  • 3.
    $\mathcal{P}(s_{t+1}|s_{t},a_{t})$: State transition distribution.
  • 4.
    $r(s_{t},a_{t})$: Reward function.
  • 5.
    $\gamma$: Discount factor.
  • 6.
    $T$: Time horizon.
In TAPF, each agent $n$ corresponds to an RL agent. The state $s_{t}$ represents the current positions of all agents. A task set Tasks contains $m$ goal positions to be assigned. Unlike MAPF, the goal position for each agent $n$ is not predefined. Instead, the problem is to formulate a goal assignment scheme (policy) for each agent. The objective is to minimize the sum of path lengths for all agents, which is equivalent to maximizing the expected cumulative rewards $\mathbb{E}_{\pi,\mathcal{P}}[r_{0:T-1}]$.
In TAPF, conflicts arise when agents’ assigned goals lead to collisions or inefficient path planning.
@@ 4.1 Modeling TAPF as a MARL Problem
As shown in Figure 1, we provide two scenarios to describe how we model the TAPF problem (described in Section 3.3) as a cooperative MARL problem. As introduced in Section 3.2, cooperative multi-agent deep RL can be modeled as a tuple $G=(\mathcal{S},\mathcal{A},\mathcal{P},r,\mathcal{O},N,\gamma,T)$. In this study, the state transition function $\mathcal{P}\left(s^{\prime}_{t+1}\mid s_{t},\boldsymbol{a_{t}}\right):\mathcal{S}\times\mathcal{A}^{N}\times\mathcal{S}\rightarrow[0,1]$ is unknown, which means the agent cannot get the environment dynamics. This is consistent with the real world TAPF problem. The number of agents $N$ in this study can be set to any integer number. The discount factor $\gamma=0.99$, which is the most common setting in deep RL. Since the main elements in multiagent deep RL are the observation space $\mathcal{O}$ (or state space $\mathcal{S}$), the action space $\mathcal{A}$ and the reward function $r$. Thus, we describe these elements in detail.
@@
Observation space: An agent’s observation contains the position and velocity itself, the relative positions of all tasks, the relative positions of the other agents, and the relative positions of neighboring obstacles. In this study, the number of other visible agents in an agent’s observation can be set to equal to or less than $N-1$. Taking Figure 1 (a) as an example, the five purple circles represent five agents, the five black circles represent five tasks. For obstacles, we set the agent can perceptive the relative positions of the adjacent obstacles, the number of adjacent obstacles can be set to any integer. In Figure 1 (a) and (b), different color curves denote the navigation trajectories of different agents.
@@ 4.2 Solving TAPF Problem via Cooperative MARL
In this study, we introduce MADDPG [31] to solve TAPF problem. Considering that agents are homogeneous, thus the agents can share the same policy network. This makes learning policy more efficiently. Let $N$ denote the number of agents, $\pi$ denote the policy parameterized by $\theta$. The centralized critic $Q$ is parameterized by $\phi$. The parameters $\theta$ for policy $\pi$ can be updated iteratively by the associated critic $Q$:
@@
where $\alpha$ is a coefficient to trade-off the weight of Q-network and target Q-network.
The share reward making the agents learn policy cooperatively. In the execution phase, we only use the policy $\pi$ whose inputs are $o_{t}^{1},\ldots,o_{t}^{N}$, respectively, and the outputs are force ($\boldsymbol{F}$) applied to the agent. According to the physical dynamics of agents described in Section 4.1, we can get the actions $\boldsymbol{a}_{t}=\left(a_{t}^{1},\ldots,a_{t}^{N}\right)$. Note that, for agent $n$, the input of $\pi$ is only the observation $o_{t}^{n}$ received by agent $n$ at time step $t$ in the execution phase.
@@ 5.1 Target Assignment and Path Planning
Figure 5 shows the training average return curves, in all the different level scenarios, the average returns increase monotonous. This verifies the stability of our method.
Figure 6 to Figure 10 show the performances target assignment and path finding in five different level scenarios. Experimental results show that although the difficulty of these five different level scenarios increasing gradually, our method performs well in all the tasks.
@@
For easy tasks (two agents - two tasks) in Figure 6 (a) (b) (c) (d), we can see that the target assignment and path finding were addressed well. For target assignment, results show that the task assignment is addressed very reasonable, because our method assigns the task to the agent that is close to the task. For path finding, it is shown that the planned paths are almost shortest.
For difficult tasks (five agents - twenty tasks) in Figure 10 (a) (b) (c) (d), we can see that the target assignment and path finding were also addressed well. For target assignment, it can be seen that each agent is reasonably assigned to several nearby tasks. For path finding, it is shown that the planned paths are usually shortest. The same superior performances can be seen in other different level scenarios, such as two agents - four tasks in Figure 7, five agents - five tasks in Figure 8, and five agents - ten tasks in Figure 9. This demonstrated the efficiency of our method.

### v2

@@ Abstract
Multi-agent target assignment and path planning (TAPF) are two key problems in intelligent warehouses. However, most literature only addresses one of these two problems separately. In this study, we propose a method to simultaneously solve target assignment and path planning from the perspective of cooperative multi-agent deep reinforcement learning (RL). This paper models the TAPF problem for intelligent warehouses to cooperative multi-agent deep RL and simultaneously addresses TAPF based on multi-agent deep RL. Furthermore, previous literature rarely considers the physical dynamics of agents. In this study, the physical dynamics of the agents are considered. Experimental results show that our method performs well in various task settings, which means that the target assignment is solved reasonably well and the planned path is almost the shortest. Moreover, our method is more time-efficient than baselines.
@@ 1 Introduction
With the rapid development of logistics-related industries, new opportunities and challenges have been proposed for intelligent warehousing [1, 2]. Traditional warehouse technologies use conveyor belts and other equipment to complete material handling. They are inflexible and not easily extensible. Intelligent warehousing utilizes a multi-agent system to deliver goods to designated locations, greatly improving warehousing efficiency. The intelligent warehouse system is divided mainly into the order fulfillment system, such as the kiva system [3] and the logistics sorting center [4]. In order fulfillment systems, mobile agents move inventory pods to inventory stations. Then, workers take goods from inventory pods, and agents move the inventory pods to their original locations. In logistics sorting centers, agents take goods from loading stations and deliver them to chutes in the center of warehouse. The approach proposed in this study is based on the logistics sorting center.
Task assignment and path finding (TAPF) are two important processes in intelligent warehousing. The system first assigns a specified task to each agent according to the order requirements. Then the agent transports the goods from the origin to the destination and ensures that the path does not conflict with other agents. The task assignment and path planning problem is typically NP-hard [5], which has a large search space. Thus, directly solving this problem is difficult. In general, two steps are required to solve the TAPF problem. The first step is multi-agent task assignment (MATA), assigning tasks to agents without considering possible path conflicts between agents. The second step is path planning for all agents by using multi-agent path finding (MAPF) algorithms. Numerous studies in the literature study MATA and MAPF; we describe these methods in detail in the next section. Although solving TAPF separately reduces the difficulty of the total problem, this type of method ignores the mutual influence between task assignment and path planning. Reasonable task assignment can not only effectively reduce the path length of warehouse agents and improve the operation efficiency, but also help avoid path conflicts between different agents. Therefore, it is necessary to solve the TAPF problems together. In the TAPF problem, we assume that each idle agent and inventory pod are homogeneous. Thus, we can assign any task to any agent.
Recently, deep reinforcement learning (RL) has received great attention due to its remarkable successes in widely challenging domains, such as robotic control [6], multi-agent problems [7], and video games [8]. In this study, we introduce cooperative multi-agent deep RL to solve the TAPF problem to address the TAPF problem simultaneously.
@@
    This paper models the TAPF problem for intelligent warehouses to cooperative multi-agent deep RL and simultaneously addresses TAPF based on multi-agent deep RL
@@
    In this paper, the physical dynamics of the agents are considered in the path planning phase.
@@
    Experimental results show that our method performs well in various of task settings, which means the target assignment is solved reasonably and the planned path is almost the shortest. Furthermore, our method is more time-efficient than the baselines.
The rest of this study is organized as follows. Section 2 discusses the related work of TAPF in an intelligent warehouse. Section 3 introduces the preliminaries of the Markov Decision Process, RL, the cooperative multi-agent deep RL, and the definition of MAPF and TAPF problem. Section 4 detailedly describes our method. Section 5 provides experimental results to verify the improved efficiency of our method. Section 6 presents conclusions and future work.
@@ 2.1 Multi-Agent Task Assignment
Through multi-agent task assignment algorithms, we can maximize the utilization of warehouse resources [9]. At present, multi-agent task assignment algorithms in intelligent warehouses can be classed into centralized and distributed classes according to the management mode. In the centralized task assignment class, a central control system is set up, which is responsible for task assignment, assigning tasks assigned to agents for execution [10]. Classical centralized task assignment algorithms include: Hungarian algorithm [11], tabu search algorithm [12], genetic algorithm [10], etc.. In the distributed assignment class, each agent in the warehouse plans its own task sequence according to tasks and environmental information [13]. This class method effectively reduces the load of the central control system and is more flexible and adaptable, but it may not find the global optimal solution [14]. Furthermore, the distributed assignment class mainly includes: learning-based methods [15] and market auction methods [16].
@@ 2.2 Multi-Agent Path Finding
In recent years, MAPF has become a hot research direction in the fields of computer science and robotics. Classical MAPF algorithms can be divided into optimal and sub-optimal types according to whether the solution results meet the optimality. The optimal typical MAPF algorithms include A*-based [17], conflict-based search [18], increasing cost tree search based [19], and compilation-based [20] methods. Sub-optimal classical MAPF algorithms include search-based [21], rule-based [22], compilation-based [23], bounded conflict-based search [24], and genetic algorithm-based [25] methods. However, traditional MAPF algorithms have poor real-time performance. Therefore, numerous researchers are beginning to study MAPF based on deep RL [26] to solve this problem. PRIMAL [27] is a typical MAPF algorithm based on deep RL. However, PRIMAL still solves the multi-agent problem by single-agent deep RL method [27]. In this study, we propose a method to simultaneously address target assignment and path planning from the perspective of cooperative multi-agent deep RL. Furthermore, in the simulation environment, the above-mentioned various MAPF algorithms rarely consider the physical dynamics of agents. In this study, the physical dynamics factors of the agent are considered.
@@ 2.3 Cooperative Multi-Agent Deep Reinforcement Learning
Cooperative multi-agent RL deals with systems consisting of several agents which interact in a common environment. Cooperative multi-agent RL aims to learn a policy that can make all agents work together to complete a common goal. Multi-agent deep RL has received great attention due to its capability to allow agents to learn to make decisions in multi-agent environments through interactions. Independent Q-learning (IQL) [28] solves the multi-agent RL problem by decomposing it into a collection of simultaneous single-agent RL problems that share the same environment. However, IQL cannot solve the nonstationarity [29] problem caused by the changing policies of other agents. Value-Decomposition Networks (VDN) [30] proposes to learn to linearly decompose a team value function into a series of per-agent value functions. After that, QMIX [7] utilizes a mixing network architecture to approximate all agents’ joint state-action value. The mixing network non-linearly combines each agent’s state-action value, which is only a condition on each agent’s local observations. Furthermore, QMIX enforces the joint state-action value monotonic with each agent state-action value by employing a network structural monotonicity constraint to the mixing network.
However, these above algorithms cannot handle continuous action space problems. The deep deterministic policy gradient (DDPG) [6] is a representative method in the continuous action task for a single agent. Multi-agent deep deterministic policy gradient (MADDPG) [31] extended DDPG from single agent setting to the multi-agent setting by using centralized training and decentralized execution paradigm [32]. MADDPG is a representative algorithm for multi-agent continuous action problems. Thus, in this study, we model the TAPF in the intelligent warehouse as a cooperative multi-agent deep RL problem and introduce MADDPG to address this problem simultaneously.
@@ 3 Preliminaries
This section provides a summary of Markov Decision Process, RL, cooperative multi-agent deep RL, and the definition of MAPF and TAPF problems.
@@ 3.3 Definition of MAPF and TAPF problem
MAPF problem: In the multi-agent path finding (MAPF) problem, the input includes an environment consisting of N agents. The environment defines all possible start and goal positions for the agents. Each agent $n$ has its own start position $start^{n}$ and goal position $goal^{n}$. In MAPF, time is typically assumed to be discrete. At time $t$, agent $n$ is at a specific location and can execute an action $a^{n}_{t}$, subsequently moving to a new position or staying at the current position. The action can be represented as $a^{n}_{t}:\text{current position}\rightarrow\text{new position}$. The path $\pi_{n}$ of agent $n$ consists of a sequence of actions $\pi_{n}=(a^{n}_{1},a^{n}_{2},\ldots,a^{n}_{T-1})$. Conflicts in MAPF typically involve position and movement collisions between agents.
TAPF problem: The TAPF problem also involves N agents within a given environment. Unlike MAPF, each agent’s goal position is not predefined. The task set includes m target positions. Each agent $n$ has a defined start position, but its goal positions are dynamically assigned based on the task set. The objective of TAPF is to define a goal position assignment scheme for each intelligent system to minimize the sum of the paths of all agents. Time is also considered discrete in this problem, with agents deciding their actions at each timestep based on their current position and assigned goals.
@@ 4.1 Modeling TAPF as a MARL Problem
As shown in Figure 1, we provide two scenarios to describe how we model the TAPF problem (described in Section 3.3) as a cooperative MARL problem. As introduced in Section 3.2, cooperative multi-agent deep RL can be modeled as a tuple $G=(\mathcal{S},\mathcal{A},\mathcal{P},r,\mathcal{O},N,\gamma,T)$. In this study, the state transition function $\mathcal{P}\left(s^{\prime}_{t+1}\mid s_{t},\boldsymbol{a_{t}}\right):\mathcal{S}\times\mathcal{A}^{N}\times\mathcal{S}\rightarrow[0,1]$ is unknown, which means the agent cannot get the environment dynamics. This is consistent with the real-world TAPF problem. The number of agents $N$ in this study can be set to any integer number. The discount factor $\gamma=0.99$ is the most common setting in deep RL. Since the main elements in multiagent deep RL are the observation space $\mathcal{O}$ (or state space $\mathcal{S}$), the action space $\mathcal{A}$, and the reward function $r$. Thus, we describe these elements in detail.
@@
Observation space: An agent’s observation contains the position and velocity itself, the relative positions of all tasks, the relative positions of the other agents, and the relative positions of neighboring obstacles. In this study, the number of other visible agents in an agent’s observation can be set to equal to or less than $N-1$. Taking Figure 1 (a) as an example, the five purple circles represent five agents, and the five black circles represent five tasks. For obstacles, we set the agent can perceive the relative positions of the adjacent obstacles, and the number of adjacent obstacles can be set to any integer. In Figure 1 (a) and (b), different color curves denote the navigation trajectories of different agents.
@@ 4.2 Solving TAPF Problem via Cooperative MARL
In this study, we introduce MADDPG [31] to solve the TAPF problem. Considering that agents are homogeneous, thus the agents can share the same policy network. This makes learning policy more efficient. Let $N$ denote the number of agents, $\pi$ denote the policy parameterized by $\theta$. The centralized critic $Q$ is parameterized by $\phi$. The parameters $\theta$ for policy $\pi$ can be updated iteratively by the associated critic $Q$:
@@
where $\alpha$ is a coefficient to trade off the weight of Q-network and target Q-network.
The shared reward makes the agents learn policy cooperatively. In the execution phase, we only use the policy $\pi$ whose inputs are $o_{t}^{1},\ldots,o_{t}^{N}$, respectively, and the outputs are force ($\boldsymbol{F}$) applied to the agent. According to the physical dynamics of agents described in Section 4.1, we can get the actions $\boldsymbol{a}_{t}=\left(a_{t}^{1},\ldots,a_{t}^{N}\right)$. Note that, for agent $n$, the input of $\pi$ is only the observation $o_{t}^{n}$ received by agent $n$ at time step $t$ in the execution phase.
@@ 5.1 Target Assignment and Path Planning
Figure 5 shows the training average return curves, in all the different level scenarios, the average returns increase monotonously. This verifies the stability of our method.
Figure 6 to Figure 10 show the performance target assignment and pathfinding in five different level scenarios. Experimental results show that although the difficulty of these five different level scenarios increases gradually, our method performs well in all the tasks.
@@
For easy tasks (two agents - two tasks) in Figure 6 (a) (b) (c) (d), we can see that the target assignment and pathfinding were addressed well. For target assignment, results show that the task assignment is addressed very reasonably because our method assigns the task to the agent that is close to the task. For pathfinding, it is shown that the planned paths are almost shortest.
For difficult tasks (five agents - twenty tasks) in Figure 10 (a) (b) (c) (d), we can see that the target assignment and pathfinding were also addressed well. For target assignment, it can be seen that each agent is reasonably assigned to several nearby tasks. For pathfinding, it is shown that the planned paths are usually the shortest. The same superior performances can be seen in other different level scenarios, such as two agents - four tasks in Figure 7, five agents - five tasks in Figure 8, and five agents - ten tasks in Figure 9. This demonstrated the efficiency of our method.